OpenStudy (anonymous):

Out of curiosity... why does this = 1?

OpenStudy (anonymous):

it doesn't

OpenStudy (anonymous):

what it the base of your logarithm?

Parth (parthkohli):

I'd be the Google. Did you mean: \(\log_5 5 = 1\)?

OpenStudy (anonymous):

@satellite73 The base is 5

OpenStudy (anonymous):

so maybe \[5^{\log_5(5)}\]?

Parth (parthkohli):

So is it \(5^{\log_5 5} = 1\)? That's 5; not 1.

OpenStudy (anonymous):

or maybe \[5^{\log_5(1)}\]

Parth (parthkohli):

Though it is \(1\) in \(\log_5 5 = 1\).

OpenStudy (anonymous):

\[a ^{\log_{a}12} = 12\]

Parth (parthkohli):

Yup :)

OpenStudy (anonymous):

So doesnt that mean \[a ^{\log_{a} } = 1 \] ?

Parth (parthkohli):

No, not really. How does the logarithm have a base but not the top part?

OpenStudy (anonymous):

Oh... right. Wow, duh haha. I don't really understand how to comprehend the ones where the log is the exponent...

Parth (parthkohli):

It's pretty easy once you get the analogy! \(\log_a b\) is another way to say, “\(a\) to the what power is \(b\)?”

Parth (parthkohli):

Let me get it a little more right.

Parth (parthkohli):

If you have two numbers \(a\) and \(b\), then \(\log _a b\) is the answer to 'what' in, “\(a\) to the what power is \(b\)?”.

OpenStudy (anonymous):

not to butt in, but this question \[a ^{\log_{a} } = 1\] doesn't make any sense log is a function, you have to take the log of something

Parth (parthkohli):

So if you ponder a bit more, you see that the thing \(a^{\log_a 12} = 12\) is just a meta-thingy.

OpenStudy (anonymous):

in other words, does not make sense to write \[\log_a\] without some variable of number after \[\log_a(x)\] or \[\log_a(75)\] or something.

Parth (parthkohli):

If you raise \(a\) to the power that you have to raise \(a\) to, for getting \(12\), you'd get 12.

Parth (parthkohli):

It's a little confusing but you'd get the point.

OpenStudy (anonymous):

@Parthkohli :O I'm going to have to think about this haha @satellite73 Yes, thank you :)

Parth (parthkohli):

@1210: You have to raise \(a\) to SOME power, to get \(12\). The "SOME" is \(\log_a 12\). Now if you raise \(a\) to "SOME" power, wouldn't you get \(12\)?

OpenStudy (anonymous):

what you need to know cold is how to switch form logarithmic form to exponential form, which really means understand what a log is \[\log_b(x)=y\iff b^y=x\] so for example \[\log_5(25)=y\iff 5^y=25\] etc

OpenStudy (anonymous):

in other words, at least as an introduction, a logarithm is an exponent if you remember that a log is an exponent, it makes all those weird rules seem rather obvious